# Tracing Multilingual Representations in LLMs with Cross-Layer Transcoders

### v1

@@ Appendix C Language Features
Appendix D Multilingual Score
We compare the top100 multilingual score with the multilingual score computed over a large training subset (Figure 14). Both entropy measures show the same overall trend across layers, with the top100 entropy consistently shifted toward lower values. The main difference appears in the earliest layers. This is likely because many features in these layers correspond to single-token activations, especially in layer 0. Given that the total number of features is limited to approximately 28K and multiple languages are involved, some features that mostly activate for a single token also tend to activate for other unrelated tokens.
@@ Appendix D Multilingual Score
We also report the general multilingual scores for Tinystories and Llama 3.2-1B in Figure 15.
Figure 15: General multilingual score for the Tinystories model and Llama 3.2-1B. The final layer jump in the Llama CLT is potentially due to training dynamics. We find that the last layer has some features activated with small values that impact slightly the reconstruction performance but have a great impact on the multilingual score.
@@
Appendix E Additional Validation Loss Curves Under Data Imbalance
In Section 4.1.1, we reported validation loss curves for four representative languages (English, Arabic, German, French). Here, we provide additional results for (i) Chinese and (ii) models trained on the TinyStories dataset.
E.1 Chinese
Figure 17 shows the validation loss curves for Chinese. The results mirror the patterns observed for other languages: despite extreme imbalance in the training data, the model maintains competitive prediction performance in Chinese, indicating that the robustness of multilingual circuits extends to typologically distinct languages.
E.2 TinyStories Models
Appendix F Additional Results: Multilingual Circuits
@@
F.1 Preposition Sentences (English: ‘‘It was a piece’’)
@@
F.2 Preposition Sentences (French: ‘‘J’ai bu une tasse’’)
@@
F.3 Content Word Sentences (English: ‘‘I prefer drinking tea to drinking’’)
@@
F.4 Content Word Sentences: ‘‘Winter, spring, summer, and autumn are the four’’
@@
F.5 Calendar Term Prediction (‘‘Monday, Tuesday, Wednesday, Thursday’’)
@@
F.6 Analogy Sentences (‘‘the opposite of ’men’ is ...’’)
@@
Appendix G Why Do Models Fail in Non-English Languages? Complementary Circuits
We provide comprehensive circuit visualizations demonstrating how circuit structures differ across languages and training mixtures, revealing the mechanisms behind performance disparities discussed in Section 4.3.
G.1 Circuit Analysis Across Mixtures
G.1.1 90% English Mixture
@@ G.2 Competing Circuit Mechanisms
Appendix H Intervention Validation
Figure 29 demonstrates successful interventions that validate the competing circuits hypothesis.
(a) 20% mixture model intervention
(b) 50% mixture model intervention
(c) 70% mixture model intervention
(d) 90% mixture model intervention
Figures  28,  27,  26, and  25 further validate our hypothesis. In particular, most of the examples where the model fails to predict the correct answer show either an absence of the reasoning circuit or a low level of its activation, and this pattern becomes less pronounced as we move towards the 20% mixture. Taken together with the intervention results, this supports our hypothesis about circuit competition.
Appendix I Category Completion Task Circuits and Interventions
Here, we present the circuits of the Category Completion Task presented in Section 4.3.2.
@@ Appendix I Category Completion Task Circuits and Interventions
Appendix J Why English Performs Better: A Case Study
@@
J.1 Cluster Activation Analysis
@@
Appendix K Latin Script Performance Advantage
Figure 34: Feature overlap between each language and English in the 20% multilingual model.
Figure 34 demonstrates that English shares more features with French and German than with Arabic and Chinese. This feature overlap explains the superior performance of Latin-script languages through shared circuit structures, even at the most balanced mixture.
Appendix L Feature Alignment with English Analysis
L.1 Methodology
We investigate whether multilingual circuits preferentially align with English representations by comparing feature alignment with native versus English tokens across layers.
For each example, we extract features contributing to the correct answer using circuit tracing. For each feature at layer $\ell$, we compute its decoder vector $\mathbf{d}_{\ell}$ and measure alignment with the unembedding space:
$\text{sim}(t)=\frac{\mathbf{d}_{\ell}}{\|\mathbf{d}_{\ell}\|}\cdot\mathbf{W}_{U}[:,t]$ (8)
where $\mathbf{W}_{U}$ is the unembedding matrix and $t$ is a token. We compute similarity for both the native answer token and its English equivalent, then rank all vocabulary tokens by similarity. Lower ranks indicate stronger alignment.
We normalize ranks to $[0,1]$ via $\text{norm}(r)=1-\frac{r-r_{\min}}{r_{\max}-r_{\min}}$, where higher values indicate better alignment. We aggregate results across layers and languages.
L.2 Results
Figure 35 shows average alignment across all languages. Both native and English tokens show similar ranking patterns across layers, with no systematic bias toward English.
(a) 20% mixture
(b) 50% mixture
(c) 70% English mixture
(d) 20% multilingual mixture
Figure 35: Feature alignment with native versus English tokens across mixture ratios. Normalized ranks show no systematic bias toward English representations across all training mixtures, with native and English tokens exhibiting similar alignment patterns throughout the network depth.

### v2

@@ D.2 Feature Overlap Results
Appendix E Multilingual Score
We compare the top100 multilingual score with the multilingual score computed over a large training subset (4M tokens)(Figure 19). Both entropy measures show the same overall trend across layers, with the top100 entropy consistently shifted toward lower values. The main difference appears in the earliest layers. This is likely because many features in these layers correspond to single-token activations, especially in layer 0. Given that the total number of features is limited to approximately 28K and multiple languages are involved, some features that mostly activate for a single token also tend to activate for other unrelated tokens.
@@ Appendix E Multilingual Score
We also report the general multilingual scores for Tinystories and Llama-3.2-1B in Figure 20.
Figure 20: General multilingual score for the Tinystories model and Llama-3.2-1B. The final layer jump in the Llama CLT is potentially due to training dynamics. We find that the last layer has some features activated with small values that impact slightly the reconstruction performance but have a great impact on the multilingual score.
@@
Appendix F Shared Multilingual Space Independence from English Dominance
To validate that the shared multilingual space structures do not require English dominance, we trained GPT-2 on balanced non-English data (25% French, 25% German, 25% Arabic, 25% Chinese) and computed weighted multilingual entropy across layers. The resulting curve exhibits the identical characteristic rise-and-fall pattern observed in English-dominant models, with peak entropy in layers 5–6 (Figure 22). This demonstrates that the shared multilngual space emergence is a fundamental architectural property, not an artifact of English dominance.
Figure 22: The Shared Multilingual Space emergence is independent of English dominance. Model trained without English (FR/DE/AR/ZH, 25% each) shows identical entropy curve to English-dominant model (20% EN + others), with characteristic rise-and-fall pattern and peak in layers 5–6.
Appendix G Case Studies: Multilingual Circuits
Throughout this section, we present circuits to validate and visualize the findings discussed in the main text. Specifically, we organize our presentation as follows:
i) Multilingual validation: We show circuits across additional tasks that support the multilingual shared space finding described in Section 4.1.3.
ii) Failure cases and interventions: We provide examples of model failures and demonstrate that interventions on circuits—primarily those that strengthen existing internal features—can successfully guide the model to produce correct outputs. This completes Section 4.3.4
iii) Translation and cultural prompts case study: We further validate our results with a case study on LLaMA-3.2-1B, using translation and culturally contextualized prompts, which confirms the robustness of our findings.
All examples produce clusters of circuits, which are manually constructed from features whose auto-interpretability descriptions and top-activating examples suggest similar functions. We further validate these clusters by intervening on them and confirming that the observed effects align with their interpreted descriptions.
G.1 Multilingual Circuits: U-shaped Entropy
@@
The top-right corner of each cluster shows its entropy score, defined as the average of the entropy values of the features it contains.
@@
G.1.1 Preposition Sentences (English: ‘‘It was a piece’’)
@@
G.1.2 Preposition Sentences (French: ‘‘J’ai bu une tasse’’)
@@
G.1.3 Content Word Sentences (English: ‘‘I prefer drinking tea to drinking’’)
@@
G.1.4 Content Word Sentences: ‘‘Winter, spring, summer, and autumn are the four’’
@@
G.1.5 Calendar Term Prediction (‘‘Monday, Tuesday, Wednesday, Thursday’’)
@@
G.1.6 Analogy Sentences (‘‘the opposite of ’men’ is ...’’)
@@
G.2 Multilingual Circuits: Failure Modes
As discussed in Section 4.3, models trained on balanced data across languages still exhibit lower performance in non-English languages compared to English. To further support the findings from Section 4.3, we examine two tasks that probe distinct types of reasoning:
  1. 1.
  2. 2.
Figure 29: Circuit structures across language mixtures. Green: 90% English mixture, orange: 50% mixture. Key clusters missing in Arabic at 90% emerge by 50%, indicating mixture-dependent circuit formation. Additional examples appear in Appendix G.2.
G.2.1 Antonym Task
Circuits driving antonym prediction are consistent across languages and mixtures, but their completeness depends on mixture balance. In Arabic under the 90% English mixture, the Men&Women cluster, present in most other languages, is missing but appears from the 50% mixture onward (Figure 29). Intervention experiments (Figure 34) confirm its causal role: adding this missing cluster restores correct Arabic predictions (Appendix G.3).
90% English Mixture
@@ G.2.1 Antonym Task
G.2.2 Category Completion Task
We analyze the Category Completion task using the 20% mixture to understand performance in the balanced setting. Arabic performs poorly due to tokenization fragmentation, which forces early layers to reconstruct words rather than activate answer-related clusters. Intervention experiments (Figure 36) show that scaling these clusters restores correct predictions, emphasizing the importance of activation strength.
We present the circuits of the Category Completion Task presented in Figures 35.
@@ G.2.2 Category Completion Task
Comparatively, English consistently triggers stronger responses in task-specific circuits (Men&Women for Antonym, Say Sports for Category Completion). This difference narrows under balanced mixtures but remains linked to sub-tokenization: languages with fragmented tokens show weaker edge strengths from embeddings to target clusters, partially explaining their reduced accuracy (Appendix I).
We further validate these mechanisms through more complex tasks (translation and cultural prompts on LLaMA-1B), tokenizer analysis, and quantitative validation across the attribution graphs (see Appendix H).
G.3 Intervention Validation
We define an intervention on a cluster of circuits as steering the residual stream using a vector constructed from the cluster’s features. Specifically, let $f_{1},f_{2},\dots,f_{n}$ denote the features in a cluster, and let $v(f_{i})$ be the decoded vector of feature $f_{i}$. The intervention vector $v_{\text{int}}$ is then given by
$v_{\text{int}}=\alpha\sum_{i=1}^{n}v(f_{i}),$ (8)
where $\alpha$ is a scaling factor determined via sweeping to optimize the effect of the intervention. The residual stream is updated by adding $v_{\text{int}}$ at the relevant layer.
Figures 34, 37 and 36 demonstrates successful interventions that validate the competing circuits hypothesis.
(a) 20% mixture model intervention
(b) 50% mixture model intervention
(c) 70% mixture model intervention
(d) 90% mixture model intervention
Figures  33,  32,  31, and  30 further validate our hypothesis. In particular, most of the examples where the model fails to predict the correct answer show either an absence of the reasoning circuit or a low level of its activation, and this pattern becomes less pronounced as we move towards the 20% mixture. Taken together with the intervention results, this supports our hypothesis about circuit competition.
Appendix H Extended Task Analysis: Translation and Cultural Prompts
We tested whether the shared multilingual space mechanisms generalize beyond simple lexical tasks (antonym, category completion) to more complex, context-dependent tasks on LLaMA-3.2-1B (Meta AI, 2026). Attribution analysis across all complex task circuits reveals the consistent three-phase structure: early layers extract language-specific features, middle layers form language-neutral semantic representations, and late layers activate language-specific output clusters.
H.1 Translation Prompts
Translation tasks require cross-lingual semantic mapping without explicit language labels:
  • “English: Flower, Français: __”
  • “English: Flower, Arabic: __”
(a) Arabic-English
(b) English-Arabic
(c) English-French
Figure 38: Circuits for the translation task.
Circuits maintain the same rise-and-fall pattern across layers, despite task complexity.
H.2 Cultural Context Prompts
Models complete: “The best dish is…” and generate language-appropriate food answers.
(a) English Best Dish
(b) German Best Dish
Figure 39: Circuits for the cultural task.
The same multilinguality pattern emerges across these circuits (see Figure 39).
Appendix I Why English Performs Better: A Case Study
@@
I.1 Cluster Activation Analysis
@@ I.2 Sub-tokenization Effects
Appendix J Model Diffing: Additional Material
J.1 Multilinguality of Repair Features
As mentionned in Section 4.3.2, we estimated the distribution of languages for features in each of our buckets. We evaluate on 10000 tokens and we report the scores in Figure 42
Figure 42: Language distribution comparison across layers and buckets for Arabic (top) and Chinese (bottom).
The blurred bars indicate features for which the language distribution could not be reliably estimated. Although we use a large number of tokens, some features are rarely activated, resulting in insufficient samples to accurately characterize their language distribution.
J.2 Token Assembly in Early Layers
Token assembly refers to the process by which early layers reconstruct sub-tokenized inputs (particularly fragmented words from morphologically rich languages like Arabic) into coherent higher-level representations before semantic processing. Logit Lens results, as described in Section 4.3.3, are shown in Figure 43. As observed, finetuning improves word assembly in Arabic and German. Counterintuitively, it increases fragmentation in Chinese, likely due to the comparatively high-rank updates required for Chinese. This suggests that the model relies more on memorization than on structured rules or heuristics for Chinese, as learning this language consumes a larger portion of the model’s capacity, leaving less capacity available for other languages.
Figure 43: Logit Lens token assembly results. We find that models tend to repair early layers by making them better at assembling tokens, thus generalizing to abstractions earlier. This especially applies to Arabic and German. Surprisingly, Chinese finetuning results in less assembly across layers. This is in accordance with the higher-rank update of Chinese that shows that the model is possibly pattern-matching.
Appendix K Tokenizer Analysis: BPE Constraints on Multilingual Processing
Despite employing a debiased multilingual BPE tokenizer trained on balanced data, morphologically rich languages exhibit severe sub-tokenization compared to other languages. We isolate the source of this disparity through three complementary analyses.
K.1 Vocabulary Allocation: Cannibalization Testing
To rule out allocation bias, we computed fertility ratios (tokens per word) for language-specific tokenizers:
  • Arabic: 1.06
  • Chinese: 1.03
  • English: 0.99
  • French: 0.96
  • German: 1.00
All ratios cluster near 1.0, confirming that the debiased tokenizer allocates vocabulary fairly across languages. The disparity in fragmentation therefore stems from linguistic properties, not allocation bias.
K.2 Morphological Coherence
Token-morpheme coherence (how well token boundaries align with morpheme boundaries) diverges dramatically:
Metric Arabic English
Token Fertility 1.97 1.53
Token-Morpheme Coherence 0.10 0.42
Meaningless Fragments (%) 89 58
Table 6: Arabic exhibits 4× lower morphological coherence than English despite fair vocabulary allocation. Meaningless fragments represent tokens that carry no semantic content.
Concrete examples illustrate this fragmentation:
  • al-radar (radar) fragments into [al-r, ad, ar]: prefix, middle, and suffix separated
  • tafaqad (inspect) fragments into [taf, qad]: broken across morpheme boundaries with subtokens having no meaning.
English maintains coherence: “reading” $\rightarrow$ [“read”, “ing”], preserving morpheme structure.
K.3 Mechanistic Implications
This tokenization constraint directly impacts early-layer processing. Fragmented tokens lack semantic content, forcing early layers (0–3) to first reconstruct word boundaries before processing meaning. This reconstruction cost weakens the initial activation patterns that propagate through multilingual circuits.
Critically, despite this early-layer handicap, Arabic still develops identical shared multilingual representations in middle layers (5–6), demonstrating that the shared multilingual space mechanism is robust to tokenization disadvantage. However, the initial signal degradation explains why Arabic exhibits weaker circuit activation and lower downstream performance despite mechanistic equivalence.